OpenStudy (anonymous):

Solve for x. 3(2x - 3) + 4(2 - x) = 0

OpenStudy (petiteme):

Use distributive property :) 6x-9+8-4x=0 add like terms 2x-1=0 transpose -1 2x=1 divide both sides by 2 x=1/2